OpenStudy (vickyisthesmartone123):

Complete the equation of the graphed linear function in point-slope form. y – 8 = (x – ) The graph of the function goes through point 2, 8 and point 6, 0.

OpenStudy (vickyisthesmartone123):

HELP PLEASE

OpenStudy (vickyisthesmartone123):

@JeruDavis @Crazy_questions

OpenStudy (jerudavis):

idk

OpenStudy (vickyisthesmartone123):

@imqwerty

OpenStudy (vickyisthesmartone123):

@jigglypuff314

OpenStudy (vickyisthesmartone123):

@mathstudent55

OpenStudy (vickyisthesmartone123):

@phi

OpenStudy (vickyisthesmartone123):

@Hero

OpenStudy (vickyisthesmartone123):

@King.Void.

OpenStudy (vickyisthesmartone123):

Help 8th grade math

OpenStudy (en):

so the first point is (2,8)?

OpenStudy (vickyisthesmartone123):

@zepdrix

OpenStudy (vickyisthesmartone123):

No ignore the bottom part and focus on the picture the bottom was nonsense

jigglypuff314 (jigglypuff314):

point slope form is \( y - k = m(x - h) \) when given that the slope is \(m\) and given a point \( (h, k)\)

OpenStudy (en):

im trying to find the slop :(

OpenStudy (en):

*slope

OpenStudy (vickyisthesmartone123):

That's the problem can someone guide me on how to find the slope?

OpenStudy (vickyisthesmartone123):

Like tell me the equations part by part and I'll answer it

jigglypuff314 (jigglypuff314):

they want us to assume that the given point for our equation should be \( (2,8)\) because they already filled in the 8 for us do you understand this? @Vickyisthesmartone123

OpenStudy (vickyisthesmartone123):

Yeah, that's basically what I only understand... :)

jigglypuff314 (jigglypuff314):

good :) now, you can figure out the slope with this idea: \(slope = \huge \frac{y_2 - y_1}{x_2 - x_1} \) if we have two points (x1, y1) and (x2, y2) which in this case can be (2,8) and (6, 0)

OpenStudy (en):

@jigglypuff314 thats it :)

OpenStudy (vickyisthesmartone123):

So...It would look like this 0-8 ----- 6-2 Yeah??? :\

jigglypuff314 (jigglypuff314):

exactly! :D now can you simplify that into a fraction? :)

OpenStudy (vickyisthesmartone123):

Umm...I think so yeah...

OpenStudy (vickyisthesmartone123):

So it would be -8 ---- 4

OpenStudy (vickyisthesmartone123):

@jigglypuff314 You there?

jigglypuff314 (jigglypuff314):

I think you can simplify that even more :)

OpenStudy (vickyisthesmartone123):

Oh ok -2 --- 1 Therefor, -2

OpenStudy (vickyisthesmartone123):

*Therefore

jigglypuff314 (jigglypuff314):

there you go, you got your slope ^_^

OpenStudy (vickyisthesmartone123):

Ok so -2 is my slope. Is that it or no?

jigglypuff314 (jigglypuff314):

correct :)

OpenStudy (vickyisthesmartone123):

That's it?

OpenStudy (vickyisthesmartone123):

I have too fill in 2 numbers

jigglypuff314 (jigglypuff314):

indeed you would also have to plug in the x value of the point (2) behind the x ;)

OpenStudy (vickyisthesmartone123):

You see

OpenStudy (vickyisthesmartone123):

-2 is my slope but whats the other part of it?

jigglypuff314 (jigglypuff314):

(2,8) if the 8 when after the y what number would go after the x in the equation? :)

OpenStudy (vickyisthesmartone123):

6?

jigglypuff314 (jigglypuff314):

look at what I said ^

OpenStudy (vickyisthesmartone123):

2? because the thing has opposite of it and on the (x-2) But on the other side it would be 2 right?

jigglypuff314 (jigglypuff314):

exactly! :D

OpenStudy (vickyisthesmartone123):

Ok so the answer is y – 8 = 2(x –2 )

jigglypuff314 (jigglypuff314):

what's the slope again?

OpenStudy (vickyisthesmartone123):

-2

jigglypuff314 (jigglypuff314):

mhmm so what's your answer again? :)

OpenStudy (vickyisthesmartone123):

y – 8 = 2(x –2 ) Or is it switched y-8=-2(x-2)

OpenStudy (vickyisthesmartone123):

Yes?

OpenStudy (vickyisthesmartone123):

:)

jigglypuff314 (jigglypuff314):

the second one is correct ^_^

OpenStudy (vickyisthesmartone123):

Ok thank you! I have something else very quick My answer was y = 0.8x + 4 am i right?

OpenStudy (vickyisthesmartone123):

@jigglypuff314 You there?

OpenStudy (vickyisthesmartone123):

JIGGLY?????!!!!! :(

jigglypuff314 (jigglypuff314):

there is no weird opposite things in the slope intercept form y = mx + b

OpenStudy (vickyisthesmartone123):

Is it correct?

jigglypuff314 (jigglypuff314):

so both the slope and the y-intercept should be negative otherwise the numbers are correct :)

OpenStudy (vickyisthesmartone123):

THANKS FOR HELPING ME @jigglypuff314 :D

OpenStudy (vickyisthesmartone123):

I already send my homework in. I submit it and got 80% WOO! Thank you! ;)